# An Approach for Auto Generation of Labeling Functions for Software Engineering Chatbots

### v1

@@ 5 Results
5.1 RQ1: How well do the generated LFs label data?
Motivation: Prior work shows that the creation of effective LFs is a tedious, time-consuming task that requires domain expertise [20]. To alleviate this burden, we propose an automated approach for generating LFs, enabling developers to focus their efforts on the core functionalities of their chatbots rather than on annotating data. Therefore, in this RQ, we evaluate the performance of our generated LFs in labeling users’ queries.
Approach: To evaluate the performance of our generated LFs, we first randomly split each studied dataset into three distinct sets: labeled, evaluation, and unlabeled data with ratios of 30%, 20%, and 50%. This approach mimics a realistic situation where unlabelled data is more abundant than labeled data, as illustrated by the motivating example in Section 2.1. Moreover, this method has been used in similar prior work [13]. We detail each data split as follows:
@@ 5.1 RQ1: How well do the generated LFs label data?
Results: Our approach generates an average number of LFs per run of 288.5 for AskGit, 19.3 for MSA, 15.7 for Ask Ubuntu, and 17.2 for Stack Overflow. The larger AskGit dataset naturally yielded more LFs, while the smaller datasets produced fewer. Table II presents the labeling performance of the generated LFs in terms of AUC and F1-score across the studied datasets. From the table, we observe that the LFs generated by our approach demonstrate promising results in labeling queries in all datasets (except Stack Overflow). Specifically, MSA, AskGit, and Ask Ubuntu exhibit AUC scores above 75%. AUC scores above 70% are generally considered good in many machine learning contexts, indicating a strong ability of the LFs to accurately differentiate between classes [28, 74]. For the Stack Overflow dataset, the generated LFs achieve an AUC score of 51.9%. To better understand the cause of the low performance on Stack Overflow, we examined the queries with mislabeled intent and found that certain intents in the Stack Overflow dataset have too few queries, making it difficult for our approach to identify distinct patterns. For example, the FacingError intent has only 10 queries in the labeled dataset, making it challenging to create an LF that can generalize to other queries for that intent.
TABLE II: Labeling Performance of the Generated LFs.
@@
Regarding the impact of using the generated LFs on the NLU’s performance, Table III presents the performance comparison between the baseline, our approach, and random labeling for the Rasa NLU in terms of AUC and F1. The difference in percentage points between our approach and the baseline is shown in ‘Approach Improvement’ column, and between our approach and random labeling in ‘Random Improvement’ column. From the table, we observe that training the NLU on data labeled by our approach improves the NLU’s performance for all datasets (except Stack Overflow). In the AskGit dataset, our approach resulted in a notable 27.2 percentage point increase in AUC. In the MSA and Ask Ubuntu datasets, the AUC increased by 8.4 and 1.9 percentage points, respectively. On the other hand, the random labeling scenario leads to a decrease in performance across all datasets, with an AUC drop of up to 43.5 percentage points. This underscores the importance of accuracy in data labeling for the NLU’s performance. In other words, increasing the training dataset size without ensuring the accuracy of the labels negatively impacts the NLU’s performance.
TABLE III: Comparison of Baseline, Our Approach, and Random Labeling Performance for Rasa NLU.
@@
5.2 RQ2: What characteristics impact the performance of the generated LFs?
Motivation: The results of RQ1 show that the LFs achieve high performance (AUC $>$ 75%) in labeling queries across most of the studied datasets. The Pruner component is employed to filter out poorly performing LFs based on specific characteristics discussed in Section 3.3. Understanding the impact of these characteristics on LF performance is crucial. This knowledge is essential not only for designing an effective pruning strategy but also for gaining a comprehensive understanding of the LFs’ overall impact on labeling performance. In this RQ, we examine the impact of these characteristics (e.g., LF’s support) on LF’s performance (i.e., AUC). Approach: To investigate the impact of LF characteristics on chatbot performance, we analyze LFs from the perspective of three key characteristics discussed in Section 3.3: Coverage, Accuracy, and LF support. For each characteristic, we group LFs based on their performance levels into high, medium, and low-performing categories. Specifically, we rank all generated LFs by each characteristic, selecting the top 20 (high), median 20 (medium), and bottom 20 (low) LFs. Next, we calculate the AUC of each labeled dataset for each characteristic group. We chose groups of 20 LFs to ensure a clear distinction between high, medium, and low-performing groups. Using larger groups diminished the clarity of these distinctions, as the performance characteristics began to overlap between categories. This approach allowed us to maintain a clear separation, ensuring the analysis accurately reflects the distinct impact of LF characteristics on performance without interference from adjacent performance levels.
TABLE IV: Labeling Performance Across Different LF Characteristics.
@@ 5.2 RQ2: What characteristics impact the performance of the generated LFs?
Results: Table IV details the impact of various LF characteristics (i.e., Coverage, Accuracy, and LF Support) on labeling performance, as measured by AUC scores across the studied datasets. Overall, we observe that all characteristics can significantly impact LF performance to varying degrees. In the following, we discuss the impact of each characteristic on the labeling performance.
Coverage: From Table IV, we observe that LFs with higher coverage consistently achieve better performance compared to those with lower coverage. For example, in the MSA dataset, LFs with high coverage achieve an AUC of 86.5% compared to 78.2% for LFs with medium coverage and 60.3& for LFs with low coverage. This performance improvement is significant for all studied datasets (except Stack Overflow), demonstrating that the proportion of data an LF can label directly impacts its performance.
Accuracy: The empirical accuracy has varying effects on labeling performance across different datasets, as shown in Table IV. For example, high-accuracy LFs achieve the best performance in the Ask Ubuntu dataset, with an AUC of 89.3%, while medium-accuracy LFs outperform high-accuracy LFs in the MSA dataset, with an AUC of 86.3% compared to 79.6%. Overall, LFs in the high and medium-performing categories outperform those in the low-performing category.
@@
6 Impact of the Number of LFs on Labeling Performance
@@
Figure 7: Impact of the number of pruned LFs on labeling performance.
7 Threats to Validity
@@
Threats to Internal Validity. Internal validity concerns factors that could have influenced our results. The choice of the threshold in our evaluation might influence the results. Using a different threshold may yield different outcomes. To mitigate this threat, we experimented with various threshold values using the MSRBot dataset [11], which is not included in our evaluation, as discussed in Section 4.3. MSRBot has a sufficient number of intents and queries and has been used in prior SE work [4, 81]. We selected the best-performing thresholds based on a manual examination of the results. Another threat is the choice of tool used for detecting entity types in the LF Generator component. To mitigate this, we relied on Rasa which has previously been shown to achieve good performance in extracting SE entities [4].
@@ 8.1 SE Chatbots
Chatbots have been developed and extensively studied across various domains, including education [82], healthcare [83], and customer service [84]. In the field of SE, they have been employed for a range of purposes where they have had a significant impact, such as assisting developers in task automation, providing guidance to newcomers, and facilitating information retrieval from software repositories [85, 86, 11, 9, 87, 88, 89].
@@
Dominic et al. [86] developed a chatbot using Rasa to assist newcomers in the onboarding process to open source projects, offering guidance, resources, and mentor recommendations. The developed chatbot helps integrate newcomers into the community more effectively. Abdellatif et al. [11] introduced MSRBot, a chatbot that answers questions extracted from software repositories, such as identifying commits that fix specific bugs. This chatbot significantly enhances the accessibility of information in software repositories. Lin et al. [9] leveraged Rasa to create MSABot, a chatbot framework aimed at supporting the development and operation of microservice-based systems, addressing challenges like modularization and scalability. Finally, Bradley et al. [88] proposed Devy, a context-aware conversational assistant that streamlines the development process by reducing manual low-level command execution, allowing developers to focus on high-level tasks.
@@
The increased attention to SE chatbots and the challenges associated with collecting data to train them [90, 86, 11] serve as the motivation for our study. Our aim is to assist practitioners in improving chatbot performance in intent classification while reducing resource costs by automating the annotation of user input. Our study differs in that we focus on supporting chatbot practitioners rather than developing chatbots ourselves.
8.2 Data Labeling
@@
There has been a substantial quantity of work on data labeling in recent years [26, 21, 91, 22]. For example, Li et al. [21] developed a method that uses a few initial rules as a starting point to identify and classify text segments, combining these rules with machine learning models. Zhao et al. introduced GLARA [91], which utilizes graph-based techniques to expand and refine rules for naming and categorizing entities in texts. Hancock et al.  [22] use natural language explanations to automatically generate LFs. This method allows users to explain their reasoning in simple language, which the system then translates into rules for data classification. Finally, Boecking et al. developed a framework [40] that enhances label generation by actively learning from user feedback. This interactive approach allows the system to continuously improve its labeling accuracy through user-guided heuristics. The work closest to ours is the work by Varma et al. [26], which proposed an approach that automates the creation of ML-based heuristics. Varma et al. evaluated their approach on medical, hardware, and text classification datasets including tasks such as image classification (bone tumor, mammogram) and sentiment analysis (Twitter sentiments).
@@ 9 Conclusion
Appendix
TABLE V: Intents, Definitions, and Query Distribution in the AskGit Dataset
Intent Number of Queries Definition

### v2

@@ 5. Results
5.1. RQ1: How well do the generated LFs label data?
Motivation: Prior work shows that manual creation of effective LFs is a tedious, time-consuming task that requires domain expertise (Galhotra et al., 2021). To alleviate this burden, we propose an automated approach for generating LFs, enabling developers to focus their efforts on the core functionalities of their chatbots rather than on annotating data. Therefore, in this RQ, we evaluate the performance of our generated LFs in labeling users’ queries.
Approach: To evaluate the performance of our generated LFs, we first randomly split each studied dataset into three distinct sets: labeled, evaluation, and unlabeled data with ratios of 30%, 20%, and 50%. This approach mimics a realistic situation where unlabelled data is more abundant than labeled data, as illustrated by the motivating example in Section 2.1. Moreover, this method has been used in similar prior work (Farhour et al., 2024). We detail each data split as follows:
@@ 5.1. RQ1: How well do the generated LFs label data?
Results: Our approach generates an average number of LFs per run of 288.5 for AskGit, 19.3 for MSA, 15.7 for Ask Ubuntu, and 17.2 for Stack Overflow. The larger AskGit dataset naturally yielded more LFs, while the smaller datasets produced fewer. Table 2 presents the labeling performance of the generated LFs in terms of AUC and F1-score across the studied datasets. From the table, we observe that the LFs generated by our approach demonstrate promising results in labeling queries in all datasets (except Stack Overflow). Specifically, MSA, AskGit, and Ask Ubuntu exhibit AUC scores above 75%. AUC scores above 70% are generally considered good in many machine learning contexts, indicating a strong ability of the LFs to accurately differentiate between classes (Yang et al., 2022; Bradley, 1997). For the Stack Overflow dataset, the generated LFs achieve an AUC score of 51.9%. To better understand the cause of the low performance on Stack Overflow, we examined the queries with mislabeled intent and found that certain intents in the Stack Overflow dataset have too few queries, making it difficult for our approach to identify distinct patterns. For example, the FacingError intent has only 10 queries in the labeled dataset, making it challenging to create an LF that can generalize to other queries for that intent.
Table 2. Labeling Performance of the Generated LFs.
@@
Regarding the impact of using the generated LFs on the NLU’s performance, Table 3 presents the performance comparison between the baseline, our approach, and random labeling for the Rasa NLU in terms of AUC and F1. The difference in percentage points between our approach and the baseline is shown in ‘Approach Improvement’ column, and between our approach and random labeling in ‘Random Improvement’ column. From the table, we observe that training the NLU on data labeled by our approach improves the NLU’s performance for all datasets (except Stack Overflow). In the AskGit dataset, our approach resulted in a notable 27.2 percentage point increase in AUC. In the MSA and Ask Ubuntu datasets, the AUC increased by 8.4 and 1.9 percentage points, respectively. On the other hand, the random labeling scenario leads to a decrease in performance across all datasets, with an AUC drop of up to 43.5 percentage points. This underscores the importance of accuracy in data labeling for the NLU’s performance. In other words, increasing the training dataset size without ensuring the accuracy of the labels negatively impacts the NLU’s performance.
Table 3. Comparison of Baseline, Our Approach, and Random Labeling Performance for Rasa NLU.
@@
5.2. RQ2: What characteristics impact the performance of the generated LFs?
Motivation: The results of RQ1 show that the LFs achieve high performance (AUC $>$ 75%) in labeling queries across most of the studied datasets. The Pruner component is employed to filter out poorly performing LFs based on specific characteristics discussed in Section 3.3. Understanding the impact of these characteristics on LF performance is crucial. This knowledge is essential not only for designing an effective pruning strategy but also for gaining a comprehensive understanding of the LFs’ overall impact on labeling performance. In this RQ, we examine the impact of these characteristics (e.g., LF’s support) on LF’s performance (i.e., AUC).
Approach: To investigate the impact of LF characteristics on chatbot performance, we analyze LFs from the perspective of three key characteristics discussed in Section 3.3: Coverage, Accuracy, and LF support. For each characteristic, we group LFs based on their performance levels into high, medium, and low-performing categories. Specifically, we rank all generated LFs by each characteristic, selecting the top 20 (high), median 20 (medium), and bottom 20 (low) LFs. Next, we calculate the AUC of each labeled dataset for each characteristic group. We chose groups of 20 LFs to ensure a clear distinction between high, medium, and low-performing groups. Using larger groups diminished the clarity of these distinctions, as the performance characteristics began to overlap between categories. This approach allowed us to maintain a clear separation, ensuring the analysis accurately reflects the distinct impact of LF characteristics on performance without interference from adjacent performance levels.
Table 4. Labeling Performance Across Different LF Characteristics.
@@ 5.2. RQ2: What characteristics impact the performance of the generated LFs?
Results: Table 4 details the impact of various LF characteristics (i.e., Coverage, Accuracy, and LF Support) on labeling performance, as measured by AUC scores across the studied datasets. Overall, we observe that all characteristics can significantly impact LF performance to varying degrees. In the following, we discuss the impact of each characteristic on the labeling performance.
Coverage: From Table 4, we observe that LFs with higher coverage consistently achieve better performance compared to those with lower coverage. For example, in the MSA dataset, LFs with high coverage achieve an AUC of 86.5% compared to 78.2% for LFs with medium coverage and 60.3& for LFs with low coverage. This performance improvement is significant for all studied datasets (except Stack Overflow), demonstrating that the proportion of data an LF can label directly impacts its performance.
Accuracy: The accuracy has varying effects on labeling performance across different datasets, as shown in Table 4. For example, high-accuracy LFs achieve the best performance in the Ask Ubuntu dataset, with an AUC of 89.3%, while medium-accuracy LFs outperform high-accuracy LFs in the MSA dataset, with an AUC of 86.3% compared to 79.6%. Overall, LFs in the high and medium-performing categories outperform those in the low-performing category.
@@
6. Discussion
Our evaluation shows that the proposed approach effectively generates LFs for labeling SE data. In this section, we discuss two important considerations for practitioners. First, we examine the relevance of NLU-based chatbots in the era of LLMs, as this context is crucial for understanding where our approach fits in the current landscape of conversational AI. Second, we analyze how the number of generated LFs affects labeling performance, since our approach can produce hundreds of LFs (e.g., 288.5 per run for AskGit) and applying all of them has computational costs. These discussions provide practical insights for deploying our approach in real-world SE chatbot development.
6.1. NLU-based Chatbots in the Era of LLMs
NLU platforms (e.g., Rasa and Dialogflow) have been the backbone of task-oriented chatbots, especially in enterprise settings. These chatbots rely on intent classifiers and entity extractors to understand user input and respond appropriately (Abdellatif et al., 2020). However, the emergence of LLMs has transformed various domains such as SE and customer services. LLMs can engage in open-ended dialogue by leveraging massive pretrained knowledge. This raises a key question: “Are traditional NLU-based chatbots still relevant for building chatbots (particularly in SE domains), or are LLMs poised to replace them?”. Recent research suggests that while LLMs dramatically expand what chatbots can do, NLU frameworks remain valuable due to their reliability, control, and efficiency (McTear et al., 2023; Foosherian et al., 2023). For example, McTear et al. (McTear et al., 2023) show that NLU pipelines still outperform LLMs on safety-critical, database-lookup tasks, while Foosherian et al. (Foosherian et al., 2023) demonstrate that a pipeline provides guard-rails that keep GPT-4’s powerful but non-deterministic outputs in check. Together, these studies underline that an NLU platform offers deterministic execution, transparent debugging, and low-latency inference, attributes that remain indispensable even as LLMs are layered on top. Below, we begin by revisiting our focus on data labeling for SE chatbots to clarify the relationship between our approach and the evolving landscape of AI models for conversational systems. We then review the technical capabilities of each approach, industry adoption trends, cost-effectiveness considerations, and other factors (like resource constraints, context handling, and efficiency) to better understand their roles in modern SE chatbots.
6.1.1. Focus on Data Labeling
Our primary contribution is an approach for automatically generating LFs that can efficiently label training data for SE chatbots. This labeled data could be used for various applications, including both training specialized NLU platforms and fine-tuning LLMs (Abdellatif et al., 2021; Liu et al., 2024). In our evaluation, we demonstrate the effectiveness of our LF-labeled data by training an NLU-based chatbot, as this provides a clearer demonstration of the impact of our labeled data compared to LLMs. This is because, unlike LLMs, which are pre-trained on vast corpora that include SE-related content (Brown et al., 2020; Liu et al., 2024), NLUs start with minimal domain-specific knowledge (Abdellatif et al., 2021). This creates a controlled environment where performance improvements can be directly attributed to our LF-labeled data. With LLMs, any performance gains might be obscured by the model’s pre-existing knowledge, making it difficult to isolate the contribution of our labeled data.
6.1.2. Why are NLU-based Chatbots still needed
Architectural Differences. Traditional NLU-based conversational chatbots use a pipeline architecture with explicit modules for NLU (intent classification and entity recognition), dialogue management (state tracking and policy), and NLG (response templates or generation) (Foosherian et al., 2023). For example, a Rasa or Dialogflow assistant will parse an input (“I lost my password”) into a structured intent (e.g., report_issue) with entities (e.g., issue_type=password), then decide the next action via defined rules or a learned policy, and finally produce a response (often from a template or small model). These NLU-based systems require developers to predefine intents, entities, and sometimes dialogue logic for the task domain (McTear et al., 2023).
Structured Control. NLU-based chatbots operate within a constrained, known action space. The dialogue manager will only execute explicitly defined actions or give responses written by chatbot developers. This makes them more predictable and easier to control (Foosherian et al., 2023). Errors are often systematic (e.g., misclassification of an intent) and can be debugged by adjusting training data or rules. As one study notes, the pipeline approach grants designers transparency and control over the agent’s behavior, helping counteract issues like hallucinations and enabling explainability for decisions (Foosherian et al., 2023). For example, a banking chatbot built with an NLU pipeline will never output an off-topic or fabricated answer—if it does not understand, it might default to a fallback message. This reliability is crucial in many SE chatbot scenarios (e.g., a DevOps assistant executing deploy commands should only do what it is programmed to do).
Reliability and Truthfulness. LLMs, while powerful, may hallucinate, generating responses that seem plausible but are actually incorrect (Foosherian et al., 2023). For example, an LLM-based chatbot might fabricate an API name or misstate a library’s capabilities when assisting a programmer, whereas an NLU-based chatbot would simply say it cannot help if it does not have that knowledge or has low confidence in its response (Abdellatif et al., 2021). It has been shown that LLM responses are not always reliable and can be factual but incorrect in subtle ways (Foosherian et al., 2023). In an SE setting, this unreliability could mean suggesting insecure code or wrong technical advice – indeed, one study noted LLMs can output “harmful advice, buggy code, or inaccurate information” without realizing it (Bocklisch et al., 2024). This is a serious concern for professional use. Additionally, LLMs lack guaranteed consistency: the same question phrased differently might yield slightly different answers (Wang et al., 2024; Zhuo et al., 2024), and they do not inherently follow a predefined style or persona (unless controlled via prompting) (Frisch and Giulianelli, 2024; Choi et al., 2025). Maintaining a consistent tone or enforcing business-specific messaging is more straightforward with template-based NLG in an NLU system, whereas with LLMs it requires careful prompt engineering or fine-tuning (Asai and Hajishirzi, 2020; Karanikolas et al., 2023; McTear et al., 2023).
Data and Learning Dependencies. While LLMs can adapt to tasks through in-context learning and fine-tuning, both methods present challenges for domain-specific applications (Dong et al., 2022). Furthermore, improving domain-specific accuracy through fine-tuning is particularly challenging due to the lack of SE benchmarks and substantial annotated datasets (Shin et al., 2023), which reinforces the need for efficient data-labeling methods like ours.
Security and Privacy. A significant consideration, especially for SE chatbots handling proprietary code or internal data, is security and privacy. Using cloud-hosted LLM APIs (like OpenAI’s) can pose risks if sensitive information is inadvertently sent to third-party servers (Ray, 2024). While enterprise tiers offer better privacy guarantees (e.g., no data retention for training) (OpenAI, 2023), many organizations, particularly in regulated sectors like finance or healthcare, prefer or require solutions that keep data entirely on-premises. Traditional NLU frameworks like Rasa can be easily deployed on-premises, offering full data control, unlike larger LLM models, which are computationally intensive to host and scale locally.
6.2. Impact of the Number of LFs on Labeling Performance
@@
Figure 7. Impact of the number of pruned LFs on labeling performance.
7. Threats to Validity
@@
Threats to Internal Validity. Internal validity concerns factors that could have influenced our results. The choice of the threshold in our evaluation might influence the results. Using a different threshold may yield different outcomes. To mitigate this threat, we experimented with various threshold values using the MSRBot dataset (Abdellatif et al., 2020), which is not included in our evaluation, as discussed in Section 4.3. MSRBot has a sufficient number of intents and queries and has been used in prior SE work (Abdellatif et al., 2021, 2024). We selected the best-performing thresholds based on a manual examination of the results. Another threat is the choice of tool used for detecting entity types in the LF Generator component. To mitigate this, we relied on Rasa which has previously been shown to achieve good performance in extracting SE entities (Abdellatif et al., 2021).
@@ 8.1. SE Chatbots
Chatbots have been developed and extensively studied across various domains, including education (et al., 2021d), healthcare (Parviainen and Rantala, 2022), and customer service (Misischia et al., 2022). In the field of SE, they have been employed for a range of purposes where they have had a significant impact, such as assisting developers in task automation, providing guidance to newcomers, and facilitating information retrieval from software repositories (Romero et al., 2020; et al., 2020c; Abdellatif et al., 2020; Lin et al., 2020; Zhang et al., 2020; Bradley et al., 2018; Serban et al., 2021).
@@
Dominic et al. (et al., 2020c) developed a chatbot using Rasa to assist newcomers in the onboarding process to open source projects, offering guidance, resources, and mentor recommendations. The developed chatbot helps integrate newcomers into the community more effectively. Abdellatif et al. (Abdellatif et al., 2020) introduced MSRBot, a chatbot that answers questions extracted from software repositories, such as identifying commits that fix specific bugs. This chatbot significantly enhances the accessibility of information in software repositories. Lin et al. (Lin et al., 2020) leveraged Rasa to create MSABot, a chatbot framework aimed at supporting the development and operation of microservice-based systems, addressing challenges like modularization and scalability. Finally, Bradley et al. (Bradley et al., 2018) proposed Devy, a context-aware conversational assistant that streamlines the development process by reducing manual low-level command execution, allowing developers to focus on high-level tasks.
@@
The increased attention to SE chatbots and the challenges associated with collecting data to train them (Santhanam et al., 2022; et al., 2020c; Abdellatif et al., 2020) serve as the motivation for our study. Our aim is to assist practitioners in improving chatbot performance in intent classification while reducing resource costs by automating the annotation of user input. Our study differs in that we focus on supporting chatbot practitioners rather than developing chatbots ourselves.
8.2. Data Labeling
@@
There has been a substantial quantity of work on data labeling in recent years (Varma and Ré, 2018; et al., 2021b; Zhao et al., 2021; et al., 2018b). For example, Li et al. (et al., 2021b) developed a method that uses a few initial rules as a starting point to identify and classify text segments, combining these rules with machine learning models. Zhao et al. introduced GLARA (Zhao et al., 2021), which utilizes graph-based techniques to expand and refine rules for naming and categorizing entities in texts. Hancock et al.  (et al., 2018b) use natural language explanations to automatically generate LFs. This method allows users to explain their reasoning in simple language, which the system then translates into rules for data classification. Finally, Boecking et al. developed a framework (Boecking et al., 2020) that enhances label generation by actively learning from user feedback. This interactive approach allows the system to continuously improve its labeling accuracy through user-guided heuristics. The work closest to ours is the work by Varma et al. (Varma and Ré, 2018), which proposed an approach that automates the creation of ML-based heuristics. Varma et al. evaluated their approach on medical, hardware, and text classification datasets including tasks such as image classification (bone tumor, mammogram) and sentiment analysis (Twitter sentiments).
@@ 9. Conclusion
Appendix A Supplementary Data
Table 5. Intents, Query Distribution and Definitions in the MSA Dataset
Intent # Queries Definition
service_api_list 11 Provides a list of APIs for a specific service.
Table 6. Intents, Query Distribution and Definitions in the Ask Ubuntu Dataset
Intent # Queries Definition
Table 7. Intents, Query Distribution and Definitions in the Stack Overflow Dataset
Intent # Queries Definition
Table 8. Intents, Query Distribution and Definitions in the AskGit Dataset
Intent # Queries Definition